# PatternKV: Flattening KV Representation Expands Quantization Headroom

### v1

@@ 4.2 Main Results
We use GSM8K to assess quantization in the non–long-text regime, adopting a zero-shot chain-of-thought paradigm. Results are shown in Fig. 5. Our method reduces accuracy loss in the non–long-text setting. This suggests that preserving the fundamental patterns of KV vectors is critical for maintaining accuracy on reasoning-intensive tasks.
@@ 4.3 Ablation Studies
Components
Table 4 shows that each component contributes positively to the overall method. Most notably, removing the adaptive threshold on the V cache leads to substantial performance degradation. This observation corroborates our earlier analysis: because semantic alignment on V varies across layers, a limited number of patterns cannot adequately cover its distribution, the nearest pattern to a given vector may thus be substantially biased, motivating a conservative rejection rule. Despite this, our approach maintains a high level of pattern utilization (about 75%). For more details, see Appendix I. We also find that leveraging patterns on K yields larger gains than on V, consistent with Hariri et al. (2025). Under low-bit settings, allocating greater quantization slack to K yields superior quantization benefits.
@@ 4.4 Efficiency and Resource Overhead Analysis
We profile inference throughput and peak memory on an NVIDIA H20 (96 GB) GPU. The input length is fixed at 1024 and the output length at 256; batch sizes are $\{16,32,48,64,96,128,160\}$. The model is Llama-3.1-8B-Instruct. As shown in Fig. 6, compared with FP16 our method attains 1.4× higher throughput and increases the single-GPU maximum batch size by 1.25×.
Figure 6: Comparison of throughput and memory footprint on Llama-3.1-8B-Instruct.
@@ 5 Related Work
Complementary to quantization, KV pruning targets redundancy by removing unimportant content before storage. Research follows two lines: (i) sequence-level token selection, where Xiao et al. (2024) retain recent tokens via sliding windows, Zhang et al. (2023); Liu et al. (2023) identify heavy hitters using attention scores, and Chitty-Venkata et al. (2025); Wang et al. (2025); Wu et al. (2024); Liu et al. (2024a); Li et al. (2024b) further improve saliency and cache stability via block eviction, one-shot top-k, soft voting, hashing, and representative snapshots; and (ii) structure-level compression, where Xu et al. (2025); Lv et al. (2025) prune low-value K/V channels and Tang et al. (2024) loads only query-relevant KV pages via query-aware metadata. Overall, historical KV caches are highly redundant, with importance concentrated in a small subset of tokens or channels.
@@ 6 Conclusion
7 Ethics Statement
All datasets used in this study are publicly available; no human subjects or annotators were involved. We confirm that our use is consistent with the datasets’ licenses and research intent, and that no personally identifiable or harmful content is included. We cite all datasets and related works accordingly.
8 Reproducibility Statement
We take several steps to ensure reproducibility: we provide detailed information on the benchmarks and their usage, the complete parameter settings for all baseline methods as well as our method, and full hardware specifications.
Appendix A Use of LLMs
We used large language models (LLMs) only as a general-purpose writing aid. LLMs did not contribute to research ideation, experiment design, implementation, analysis, or result interpretation, and no text was directly copied without human review. No proprietary or sensitive data were provided to LLMs. All technical content, claims, and conclusions are authored and verified by the authors.
Appendix B K Cache Pattern Stable Analysis
@@
Appendix C Supplementary Figures For Insight 1 And Insight 2
We provide additional experimental observations that corroborate our insights. See Figs. 7 and 8.
@@ Appendix C Supplementary Figures For Insight 1 And Insight 2
Appendix D Additional Proof
Goal. Show that for any bit-width $b\geq 1$ and any $\rho\in(0,1)$, there exists a finite pattern set $\mathcal{P}$ such that the residual scheme attains a uniform worst-case error bound satisfying
@@ Appendix D Additional Proof
Appendix E Detailed Information of LongBench
@@
Appendix F Baseline Settings
@@
Appendix G INT4 Results on LongBench
@@
Appendix H INT4 Results on Long-CoT settings
@@
Appendix I V Pattern Utilization Rate

### v2

@@ 4.2 Main Results
We use GSM8K to assess quantization in the non–long-text regime, adopting a zero-shot chain-of-thought paradigm. All results in Fig. 5 are obtained under the INT2 quantization setting. Our method reduces accuracy loss in the non–long-text setting. This suggests that preserving the fundamental patterns of KV vectors is critical for maintaining accuracy on reasoning-intensive tasks.
Figure 5: GSM8K accuracy under zero-shot CoT on Llama-3.1-8B-Instruct with an INT2 quantization setting.
@@ 4.3 Ablation Studies
Components
Table 3 shows that each component contributes positively to the overall method. Most notably, removing the adaptive threshold on the V cache leads to substantial performance degradation. This observation corroborates our earlier analysis: because semantic alignment on V varies across layers, a limited number of patterns cannot adequately cover its distribution, the nearest pattern to a given vector may thus be substantially biased, motivating a conservative rejection rule. Despite this, our approach maintains a high level of pattern utilization (about 75%). For more details, see Appendix H. We also find that leveraging patterns on K yields larger gains than on V, consistent with (Hariri et al., 2025). Under low-bit settings, allocating greater quantization slack to K yields superior quantization benefits.
@@ 4.4 Efficiency and Resource Overhead Analysis
PatternKV is designed to further reduce quantization error, introducing only modest system overhead compared to KIVI. The primary source of overhead is a one-time pattern-mining step in prefill, and even this single step adds only 6% to latency. Despite this, our method delivers 1.5× higher throughput than FP16 and increases the maximum single GPU batch size by 1.25×. While KIVI employs a simpler, highly optimized pipeline which limits its accuracy ceiling, PatternKV delivers significant accuracy improvements with only a minimal reduction in speed. Moreover, PatternKV consistently outperforms heavier baselines such as ZipCache and SKVQ in both latency and throughput, and remains competitive with OTT while providing clear accuracy improvements. For memory overhead, PatternKV is near lossless in practice, increasing peak GPU memory by only about 0.42% compared with KIVI. In contrast, several alternative approaches incur noticeably higher peak-memory pressure, whereas PatternKV keeps the peak GPU memory footprint essentially on par with KIVI while improving generation quality. A more detailed efficiency analysis is provided in Appendix I.
@@ 5 Related Work
Complementary to quantization, KV pruning targets redundancy by removing unimportant content before storage. Research follows two lines: (i) sequence-level token selection, where Xiao et al. (2024) retain recent tokens via sliding windows, Zhang et al. (2023); Liu et al. (2023) identify heavy hitters using attention scores, and Chitty-Venkata et al. (2025); Wang et al. (2025); Wu et al. (2024); Liu et al. (2024a); Li et al. (2024b) further improve saliency and cache stability via block eviction, one-shot top-k, soft voting, hashing, and representative snapshots; and (ii) structure-level compression, where Xu et al. (2025); Lv et al. (2025) prune low-value KV channels and Tang et al. (2024) loads only query-relevant KV pages via query-aware metadata. Overall, historical KV caches are highly redundant, with importance concentrated in a small subset of tokens or channels.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A K Cache Pattern Stable Analysis
@@
Appendix B Supplementary Figures For Insight 1 And Insight 2
Settings
All preliminary experiments were conducted on the GSM8K dataset. To ensure fair comparison, we employed a unified zero-shot chain-of-thought (CoT) prompt template across all models and experimental conditions. This design eliminates variability due to prompt formatting and allows a cleaner assessment of the impact of the proposed method.
[…]
We provide additional experimental observations that corroborate our insights. See Figs. 6 and 7.
@@ Appendix B Supplementary Figures For Insight 1 And Insight 2
Appendix C Additional Proof
Goal Show that for any bit-width $b\geq 1$ and any $\rho\in(0,1)$, there exists a finite pattern set $\mathcal{P}$ such that the residual scheme attains a uniform worst-case error bound satisfying
@@ Appendix C Additional Proof
Appendix D Detailed Information of LongBench
@@
Appendix E Baseline Settings
@@
A potential source of confusion is the fact that the residual size differs between KIVI and OTT. This discrepancy does not bias the comparison; on the contrary, it is introduced to make the comparison fairer. In KIVI, the quantization strategy on the K cache does not maintain a continuous sliding window: after each group is quantized, that group is immediately cleared from the residual buffer. In contrast, OTT maintains a fixed sliding window of length residual_size throughout the entire decoding process, so that the last residual_size tokens always remain in higher precision.
To keep the effective number of FP16 tokens preserved during decoding aligned across methods, we therefore set OTT’s $\text{residual\_size}=32$. This choice ensures that the number of unquantized tokens retained by OTT at each decoding step is comparable, on average, to the number of FP16 tokens maintained by KIVI under its group-based quantization behavior.
Appendix F INT4 Results on LongBench
@@
Appendix G INT4 Results on Long-CoT settings
@@
Appendix H V Pattern Utilization Rate
@@
Appendix I Detailed Efficiency Analysis of PatternKV
In this section, we provide a component-level breakdown of the computational overhead introduced by PatternKV, together with head-to-head comparisons against KIVI, ZipCache, OTT, and SKVQ in terms of latency, throughput, and peak GPU memory.
Settings
We profile inference throughput and peak memory usage on an NVIDIA H20 (96 GB) GPU. The input and output lengths are fixed at 1024 and 256 tokens, respectively, and we evaluate batch sizes of $\{16,32,48,64,96,128,160\}$. The model is Llama-3.1-8B-Instruct.
I.1 Component-Level Breakdown
The additional computational components of PatternKV, beyond standard KV quantization, are:
  • Pattern mining (prefill). K-means clustering over $|\mathcal{M}|$ patterns during the prefill stage.
  • Chebyshev-center updates (decode). Non-iterative updates of pattern centers as decoding proceeds.
  • Pattern selection (decode). At each decoding step, selecting the appropriate pattern via an index lookup and gathering the corresponding centroids.
Table 9: Component-level latency, throughput, and FLOPs associated with the PatternKV-specific procedures (batch size =32=32). Latency and throughput are reported as relative change with respect to the baseline attention pipeline (negative throughput indicates a drop).
Component Latency Throughput GFLOPs
Pattern mining $6.60\%$ $6.10\%$ 6.91
Pattern selection $2.59\%$ $2.47\%$ 2.07
Chebyshev-center updates $3.11\%$ $2.94\%$ 2.78
Table 9 reports the overhead of three PatternKV-specific procedures at batch size $32$. The dominant source of overhead is the pattern mining step in the prefill stage, which is expected since K-means requires iterative updates to converge. In comparison, pattern selection and Chebyshev-center updates each contribute only about 2%-3% additional latency. Overall, these components introduce only modest extra cost relative to the full attention pipeline, while enabling the accuracy gains brought by PatternKV.
I.2 Head-to-Head Latency and Throughput
Under identical hardware, model, quantization configuration, batch sizes, and sequence lengths, we compare end-to-end generation latency and throughput across FP16, KIVI, OTT, ZipCache, SKVQ, and PatternKV.
Latency
Table 10 summarizes the per-token latency (in seconds) for different batch sizes.
Table 10: End-to-end latency (s) under different batch sizes. “–” indicates that the corresponding configuration does not fit in GPU memory.
Method bz=16=16 bz=32=32 bz=48=48 bz=64=64 bz=96=96 bz=128=128 bz=160=160
FP16 11.29 18.39 24.46 30.57 45.78 58.70
KIVI 11.48 12.10 13.37 15.13 22.29 27.30 36.70
OTT 10.34 12.72 15.23 17.59 25.93 31.96
ZipCache 23.02 40.84 57.27 74.04 109.89 143.13
SKVQ 16.73 26.87 77.95 182.45 220.83
PatternKV 9.80 13.49 16.73 19.86 29.36 36.69 48.37
Throughput
Table 11 reports the corresponding throughput (tokens per second).
Table 11: End-to-end throughput (tokens/s) under different batch sizes. “–” indicates that the corresponding configuration does not fit in GPU memory.
Method bz=16=16 bz=32=32 bz=48=48 bz=64=64 bz=96=96 bz=128=128 bz=160=160
FP16 1089.33 1337.54 1508.60 1609.53 1612.52 1676.64
KIVI 1071.59 2032.88 2759.44 3251.66 3311.32 3605.49 3352.05
OTT 1189.93 1934.07 2423.10 2796.90 2846.63 3079.22
ZipCache 534.36 602.51 644.50 664.68 671.78 687.67
SKVQ 735.25 915.52 473.48 269.74 334.29
PatternKV 1255.05 1823.69 2206.27 2477.12 2514.12 2682.75 2543.68
From Tables 10 and 11, we observe that PatternKV incurs an increase in latency and a decrease in throughput compared with KIVI at larger batch sizes. This is expected, as PatternKV introduces additional components into the inference pipeline. At the same time, PatternKV delivers substantial accuracy improvements over KIVI (see main text), leading to a more favorable overall trade-off between efficiency and performance. Compared with ZipCache and SKVQ, PatternKV attains consistently better latency and throughput across all reported batch sizes, and it is competitive with OTT while being more accurate.
I.3 Peak Memory Overhead
We next provide a fine-grained analysis of the memory overhead of PatternKV and validate it using empirical peak GPU memory measurements.
Analytical memory estimates
We maintain a separate set of centroids for each attention head, which introduces additional memory overhead. Let $L$ denote the number of layers, $H$ the number of KV heads, $B$ the batch size, and $S$ the sequence length (in tokens). For INT2 quantization, each stored value occupies $2$ bytes.
Prefill stage In the prefill stage, the number of pattern vectors is fixed regardless of the context length. The memory footprint of pattern vectors can be approximated as
$M_{\mathrm{pattern,prefill}}=2LH\cdot 32\cdot 2\ \mathrm{bytes}$
where the leading factor $2$ accounts for both $K$ and $V$, and $32$ is the number of pattern centroids per head. The KV cache memory usage is
$M_{\mathrm{KV}}=2BLHS\cdot 2\ \mathrm{bytes}$
The relative memory fraction of pattern vectors during prefill is therefore
$\frac{M_{\mathrm{pattern,prefill}}}{M_{\mathrm{KV}}}=\frac{32}{BS}$
For a medium-length sequence with $S=8192$ (approximately $8\mathrm{K}$) and $B=1$, the pattern vectors account for only about $0.39\%$ of the KV memory. As the batch size increases, this fraction decreases further.
Decode stage In the decode stage, we set $G_{\mathrm{pattern}}=128$, meaning that one pattern vector is generated for every $128$ decoding steps. Analogously, the relative memory ratio of pattern vectors during decoding is
$\frac{M_{\mathrm{pattern,decode}}}{M_{\mathrm{KV}}}=\frac{1}{BG_{\mathrm{pattern}}}$
This ratio is at most $0.78\%$ when $B=1$, and again decreases as the batch size grows. In our implementation, the extra centroid tensors account for only about $0.42\%$ of the overall KV-cache memory footprint.
Empirical peak GPU memory
Table 12 reports the peak GPU memory of each method during inference, which corroborates the above estimates.
Table 12: Peak GPU memory consumption (GB) under different batch sizes.
Method bz=16=16 bz=32=32 bz=48=48 bz=64=64 bz=96=96 bz=128=128 bz=160=160
FP16 21.93 28.84 35.83 42.77 56.67 70.56
KIVI 21.19 27.39 33.57 39.77 52.16 64.55 76.95
OTT 24.07 31.13 38.15 45.20 59.31 73.41
ZipCache 23.10 29.21 35.35 41.43 53.61 65.84
SKVQ 24.68 32.37 40.06 47.75 63.13
PatternKV 21.23 27.48 33.69 39.90 52.39 64.81 77.28
Compared with the KIVI baseline, PatternKV introduces only about $0.42\%$ additional peak memory overhead while yielding a clear improvement in performance. This indicates that a very small extra storage cost is sufficient to obtain meaningful performance gains with PatternKV. In contrast, OTT, ZipCache, and SKVQ exhibit higher peak memory usage and less favorable scaling with the batch dimension, largely due to limitations in the publicly available implementations. Both KIVI and PatternKV maintain a well-controlled peak memory profile across the evaluated batch sizes.
Appendix J K-drift under ALiBi positional encoding
To further examine whether the observed drift of K patterns is primarily induced by RoPE, we additionally conduct pattern-evolution experiments on MPT-7B-Chat, a decoder-only model that adopts ALiBi positional encoding. Unlike RoPE-based models, the K cache in MPT-7B-Chat does not exhibit a clear progressive evolution with sequence length: tokens from early and late positions can share similar patterns, and abrupt changes may occur even between adjacent positions. Figure 9 visualizes the pattern assignments along the sequence dimension for representative attention heads. As shown in the figure, the pattern assignments under ALiBi remain relatively mixed across positions, without the smooth positional drift observed in the RoPE-based setting. These observations provide further evidence that the systematic drift of K patterns reported in the main paper is largely a consequence of RoPE, rather than an inherent property of attention weights or of the proposed pattern-mining procedure.
Figure 9: Pattern evolution of the K cache under ALiBi positional encoding on MPT-7B-Chat. Each row corresponds to an attention head, and colors denote discovered K patterns. Tokens from early and late positions can share similar patterns and abrupt pattern changes may occur between adjacent positions, in contrast to the smooth K-drift observed in RoPE-based models (see Figure 3).
Appendix K PatternKV under Multi-Head Latent Attention (MLA)
To assess how PatternKV extends to models with Multi-Head Latent Attention (MLA) and other non-standard attention architectures, we take DeepSeek-V2-Lite-Chat as a representative MLA model. Since, in MLA, RoPE is no longer applied to the entire K cache, we first verify whether the K cache in MLA still exhibits the key properties observed in GQA/MHA models.
Based on Figure 10 , we make the following observations:
  • The RoPE-applied part of the K cache in MLA has substantially larger magnitudes than the non-RoPE part, which further suggests that RoPE is the primary source of the outlier distribution in the K cache.
  • The MLA K cache remains stable along the channel dimension, indicating that a small number of pattern vectors can still effectively cover its distribution.
  • Along the sequence dimension, MLA exhibits the same gradual evolution behavior as GQA/MHA, which can be attributed to the RoPE information injected into the cache.
These findings indicate that PatternKV can be directly applied to the KV cache that is actually stored in MLA, without requiring any additional architectural adaptations. In future work, we plan to conduct more systematic MLA-based experiments to further validate these observations and to explore broader design choices for pattern-based KV compression in non-standard attention architectures.
Figure 10: Pattern evolution of the K cache in DeepSeek-V2-Lite-Chat (MLA). The RoPE-applied subspace exhibits larger magnitudes and a gradual evolution along the sequence dimension, while remaining stable along channels, which mirrors the behavior observed in GQA/MHA and supports the applicability of PatternKV to MLA.